A perfectly competitive fir has the following cost function:

1/3q3-9/2π‘ž2 +10π‘ž+500 32

and demand and supply functions are also given as:

𝑄𝐷 =50βˆ’3𝑃 and 𝑄𝑆 =10+2𝑃

What is the equilibrium price and quantity in this market?

What is the profit maximizing quantity that this firm would produce?


1
Expert's answer
2021-06-22T14:05:32-0400

"QD = 50-3P"

"QS= 10 + 2P"

In finding Equilibrium price,Β 

"QS = QD"

"50 -3P= 10+2P"

"50-10 = 2P + 3P"

"40 = 5P"

"P = 40\/5"

"P = 8"

ii) "Equilibrium quantity"

In finding Equilibrium quantity we use the QD equation

"QD = 50-3P"

"= 50-3(8)"

"= 50-24"

"Equilibrium quantity = 26"

2). Profit Maximizing quantity

"P = MC"

"P = 8"

"MC = dTC\/dQ"

"1\/3q^3 -9\/2q^2 +10q +50032"

"3\/3q^2 +18\/2q+10"

"= 1\/q^2 + 9\/q +10"

"MC = 2\/q+ 9"

"= 2\/q + 9"

"8 = 2\/q+ 9"

"2\/q= 9-8"

"2\/q = 1"

"q = 2"

"Profit Maximizing quantity = 2"